OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 10 cubic meters. The length of its base is twice the width. Material for the base costs 14 dollars per square meter. Material for the sides costs 8 dollars per square meter. Find the cost of materials for the cheapest such container.

OpenStudy (anonymous):

Volume: V = w L h <=> 10 = 2 w² h -> h = 5/ w² Surface Area: bottom + sides SA = 2w² + 2 wh + 2 ( 2w h) Replace h = 5/w²: = 2( w² + 5/w + 10/w) = 2 ( w² + 15/w ) Cost C = 2 ( 14 w² + 120/w ) Minimize Cost, C' = 0 C' = .... Thus C (w) = ... @sabadoosh Questions?